OpenStudy (anonymous):

what is the equation of the line perpendicular to the line y=4/9x-2 that passes through the point (4,3)

OpenStudy (anonymous):

Need some clarification. Is this (4/9)x-2, (4/9x)-2, or 4/(9x-2)?

OpenStudy (anonymous):



OpenStudy (anonymous):

slope will be=-9/4 equation, (y-3)=-9/4(x-4)

OpenStudy (nirmalnema):

the eq u will get after simplifying it is 9y=4x-18use one point form (y-3)= 4/9 (x-4) 9y-27=4x-16 so u will get 4x-9y +11

OpenStudy (anonymous):

@nirmalnema a. y=-4/9x-2 b. y=9/4x+6 c. y=-9/4x+12 d. y=-9/4x-12

OpenStudy (nirmalnema):

@cutegirl what are these..

OpenStudy (anonymous):

thats the answer i have to choose from

OpenStudy (nirmalnema):

sorry i told u wrong ans by mistake i told u the ans of parallel line...really sorry

OpenStudy (anonymous):

its okay :)

OpenStudy (nirmalnema):

then slope will be=-9/4 and equation ,is (y-3)=-9/4(x-4) 4y-12=-9x+36 4y=-9x+48 y=-9/4x + 12

OpenStudy (anonymous):

thanks so much :)))

OpenStudy (nirmalnema):

@cutegirl ur welcome